### v2

@@ V.1 Bosonic Abelian phase
It was already shown in Ref. Kobayashi et al., 2024a that $\{\zeta_{n}\}$ can all be extracted from the partial rotation. In addition, both $\sigma_{H}$ and $c_{-}$ are shown to be extracted from partial rotation and topological entanglement entropy in Sec. III.2.1. These span all the obstructions to $\mathrm{U}(1)$-preserving gapped edge state mentioned above.
@@ VI Lieb-Schultz-Mattis type constraint
Suppose that we observed $\{\mathcal{T}_{1}(2\pi/p_{j},\pi/p_{j})\}$ has a non-trivial phase for a set of distinct prime numbers $\{p_{j}\}$. One can see that this leaves us two possibilities: 1. the edge cannot be gapped while preserving $\mathrm{U}(1)^{f}$ symmetry, or 2. the edge is gappable preserving $\mathrm{U}(1)^{f}$ symmetry, where the bulk must have multiple ground states on a torus with the degeneracy non-trivially bounded from below. This statement is reminiscent of the Lieb-Schultz-Mattis type theorems Lieb et al. (1961); Oshikawa (2000); Hastings (2004), which constrains the low-energy spectrum for a given input of the symmetry action on the ground state. Concretely, one can show the following theorem:
Theorem. For a (2+1)D topological order with $\mathrm{U}(1)^{f}$ symmetry, suppose that the partial rotation $\{\mathcal{T}_{1}(2\pi/p_{j},\pi/p_{j})\}$ has a non-trivial phase for a set of distinct prime numbers $\{p_{j}\}$. Then, the topological order admits a $\mathrm{U}(1)^{f}$ symmetry-preserving gapped edge state only if the number of anyons $r$ satisfies $r\geq r_{0}$, where $r_{0}$ is the smallest integer satisfying $\prod_{j}p_{j}\leq 2^{\frac{4}{3}r_{0}+8}3^{\frac{4}{3}r_{0}}$.
Proof. Let us assume that the topological order admits a $\mathrm{U}(1)^{f}$-preserving gapped edge, while the phases of $\{\mathcal{T}_{1}(2\pi/p_{j})\}$ are non-trivial. $\mathcal{T}_{1}(2\pi/p_{j},0)$ obstructs the $\mathrm{U}(1)^{f}$-preserving gapped edge when $\gcd({N_{\mathrm{FS}}}(\mathcal{C}),p_{j})=1$, where $\mathcal{C}$ is the super-modular tensor category that describes the anyons of the topological order. We hence must have $\gcd({N_{\mathrm{FS}}}(\mathcal{C}),p_{j})>1$. This implies that ${N_{\mathrm{FS}}}(\mathcal{C})$ must be divisible by $\prod_{j}p_{j}$, so we have $\prod_{j}p_{j}\leq{N_{\mathrm{FS}}}(\mathcal{C})$. Meanwhile, let us consider a modular tensor category $\breve{\mathcal{C}}$ obtained by gauging ${\mathbb{Z}}_{2}^{f}$ symmetry of the fermionic topological phase $\mathcal{C}$. Mathematically, $\breve{\mathcal{C}}$ is the minimal modular extension of the super-modular category $\mathcal{C}$. For a modular category, it is known that ${N_{\mathrm{FS}}}(\breve{\mathcal{C}})\leq 2^{\frac{2\breve{r}}{3}+8}3^{\frac{2}{3}\breve{r}}$, where $\breve{r}$ is the number of anyons in $\breve{\mathcal{C}}$ Bruillard et al. (2015). In the presence of $\mathrm{U}(1)^{f}$ symmetry, there is an Abelian vison $v$ that carries ${\mathbb{Z}}_{2}^{f}$ vortex. In that case, one can write $\breve{r}=2r$, where $r$ is the number of anyons in $\mathcal{C}$. Using ${N_{\mathrm{FS}}}(\mathcal{C})\leq{N_{\mathrm{FS}}}(\breve{\mathcal{C}})$, we have
@@ VII.2 Review: anomaly and Hall conductivity of topological order with Lie group symmetry
$\begin{split}&\sum_{j}2\pi h(\pi_{\mathcal{A}}({\bf k}_{j}))\int_{M^{4}}\mathcal{P}(\pi_{\mathcal{A}}(w_{2}^{j}))\\ &+\sum_{j<k}2\pi m(\pi_{\mathcal{A}}({\bf k}_{j}),\pi_{\mathcal{A}}({\bf k}_{k}))\int_{M^{4}}\pi_{\mathcal{A}}(w_{2}^{j})\cup\pi_{\mathcal{A}}(w_{2}^{k})~{},\end{split}$ (39)
@@
$\sum_{j}\int_{C}\pi_{f}(g_{b}^{\ast}w_{2}^{j})=\int_{C}w_{2}(TM^{4})\quad\bmod 2,$ (40)
for any 2-cycle $C$. $w_{2}(TM^{4})$ is the second Stiefel-Whitney class of the tangent bundle of the spacetime $M^{4}$ on which the response action is supported.
@@ VII.4 Example: $G_{f}=\mathrm{SU}(2)^{f}$
$\frac{2\pi k}{4}\int_{M^{4}}\mathcal{P}(w_{2}(\mathrm{SO}(3))).$ (43)
This response action of $\mathrm{SU}(2)^{f}$ symmetry actually corresponds to a trivial SPT, when the Dirac quantization condition (40) is understood; $w_{2}(\mathrm{SO}(3))=w_{2}(TM^{4})$. In that case we have $\int\mathcal{P}(w_{2}(TM^{4}))=\sigma(M^{4})$ mod 4, with $\sigma(M^{4})$ the signature of a 4-manifold $M^{4}$. Hence, the above response action (43) is identified as the continuous gravitational theta angle, and topological order does not carry ’t Hooft anomaly of SU(2)f symmetry.
@@ VIII Discussions
Finally, the partial operations also allow us to characterize the invariants of crystalline topological phases with point group symmetry Shiozaki et al. (2017); Zhang et al. (2023); Kobayashi et al. (2024b). It would be interesting to develop a theory for the crystalline invariants of topological order, which is relevant to various invariants characterizing the crystalline spin liquids.
@@ A.1 Fermionic case
in terms of edge CFT with $\mathrm{U}(1)^{f}$ symmetry by utilizing the cut-and-glue approach Qi et al. (2012). For simplicity, let us consider the case where the edge CFT is described in terms of chiral rational CFT; the generalization to non-chiral case is rather straightforward, see Ref. Kobayashi et al., 2024a for the discussions in bosonic non-chiral phases.
@@
where we introduced the velocity $v$, correlation length $\xi_{l}=v\beta_{l}$, $\xi_{r}=v\beta_{r}$, and the circumference of the cylinder $L$. ${P}_{l}$ and ${P}_{r}$ are translation operators on the left and right edge ${P}_{l}=-\frac{1}{v}H_{l}$, ${P}_{r}=\frac{1}{v}H_{r}$. $\chi_{\alpha}(\tau;[s,\theta],[s^{\prime},\theta])$ with $s,s^{\prime}\in\{\mathrm{AP},\mathrm{P}\}$, $\theta,\theta^{\prime}\in[0,\pi)$ is the CFT character that corresponds to the partition function on a torus with modular parameter $\tau$ equipped with Spinc structure. For example,
@@
where $\alpha$ labels the topological sector within the anti-periodic sector, with the $\mathrm{U}(1)$ twisted boundary condition by $\theta$ in the spatial cycle. In our setup where $L\ll\xi_{l}$, the CFT characters for the left edge are approximated as
@@ A.1.1 The case where $n$ is odd
$\begin{split}\chi_{a}\left(\frac{i\xi_{r}}{L}-\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}]\right)&=S_{ab}\chi_{b}\left(-\frac{1}{\frac{i\xi_{r}}{L}-\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\\ &=(ST^{n})_{ab}\chi_{b}\left(\frac{-in\frac{\xi_{r}}{L}}{\frac{i\xi_{r}}{L}+\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\end{split}$ (49)
@@
Here, $[]_{\pi}$ denotes the mod $\pi$ operation and the spin structure is acted by ${\mathbb{Z}}_{2}$ action as $\mathrm{AP}+1=\mathrm{P}$, $\mathrm{P}+1=\mathrm{AP}$. To evaluate the action of $S,T$ modular matrices, we note that the modular transformation in the twisted Hilbert space of CFT can be read by $S,T$ matrices of $G$-crossed braided fusion category that describes the topological order with global symmetry in the bulk Barkeshli et al. (2019). In our case, the bosonic symmetry group is $G_{b}=\mathbb{R}/(\pi\mathbb{Z})$, and the relevant subgroup for the operation $U_{\frac{\pi}{n}}$ is the ${\mathbb{Z}}_{n}$ subgroup of $G_{b}$. Using the data of ${\mathbb{Z}}_{n}$-crossed braided fusion category, the $T$-matrix element is expressed as Barkeshli et al. (2019)
@@
$\begin{split}\chi_{a}\left(\frac{i\xi_{r}}{L}-\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}]\right)&=\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}(ST^{n})_{ab}\chi_{b}\left(\frac{-in\frac{\xi_{r}}{L}}{\frac{i\xi_{r}}{L}+\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\\ &=e^{-\frac{2\pi in}{24}c_{-}}\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}S_{ab}\theta_{b}^{n}\prod_{j=0}^{n-1}\eta_{b}\left(\frac{\pi}{n},\frac{j\pi}{n}\right)\cdot\chi_{b}\left(\frac{-in\frac{\xi_{r}}{L}}{\frac{i\xi_{r}}{L}+\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\\ &=e^{-\frac{2\pi in}{24}c_{-}}\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}\sum_{c\in\mathcal{C}_{\mathrm{AP},0}}S_{ab}\theta_{b}^{n}\prod_{j=0}^{n-1}\eta_{b}\left(\frac{\pi}{n},\frac{j\pi}{n}\right)\cdot S_{bc}\chi_{c}\left(\frac{iL}{n^{2}\xi_{r}}+\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}]\right)\end{split}$ (52)
@@
$\mathcal{T}^{f}_{a}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)=e^{\frac{2\pi i}{n}(h_{a}-\frac{c_{-}}{24})}\frac{e^{-\frac{2\pi in}{24}c_{-}}\sum_{b,c}(S_{a,b}+S_{a\psi,b})\theta_{b}^{n}\prod_{j=0}^{n-1}\eta_{b}\left(\frac{\pi}{n},\frac{j\pi}{n}\right)\cdot S_{bc}\chi_{c}(\frac{iL}{n^{2}\xi_{r}}+\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}])}{\sum_{b}(S_{a,b}+S_{a\psi,b})\chi_{b}(\frac{iL}{\xi_{r}};[\mathrm{AP},0],[\mathrm{AP},0])}$ (53)
@@
where $\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}$ is the twisted sector in $G_{b}$-crossed extension of super-modular category that consists of anyons in the $[\mathrm{AP},\frac{\pi}{n}]$ sector. To relate this quantity to $\zeta^{H}_{n}$, we express the formula in terms of the anyons in the untwisted sector, instead of the twisted sector by $\frac{\pi}{n}$. This is done by using the consistency equations and the Verlinde formula of $G$-crossed braided fusion category given by Barkeshli et al. (2019)
@@
where we labeled ${\bf g}=\pi z_{1},{\bf h}=\pi z_{2},{\bf k}=\pi z_{3}$ with $z_{1},z_{2},z_{3}\in[0,1)$, and $v\in\mathcal{C}_{\mathrm{P}}$ is the vison defined as the relation $M_{b,v}=e^{i\pi Q_{b}}$ for all anyons $b\in\mathcal{C}_{0},\mathcal{C}_{\mathrm{P}}$. The parameter $k\in{\mathbb{Z}}$ corresponds to stacking the bosonic IQH phase atop the state (here, the bosonic IQH is equivalent to eight copies of fermionic IQH state stacked with $\overline{E}_{8}$ state, so carries $\sigma_{H}=8$). The combination $h_{v}+k$ is identified with the electric Hall conductivity of the bosonic topological order as Manjunath and Barkeshli (2020)

### v3

@@ V.1 Bosonic Abelian phase
It was already shown in Ref. Kobayashi et al., 2024b that $\{\zeta_{n}\}$ can all be extracted from the partial rotation. In addition, both $\sigma_{H}$ and $c_{-}$ are shown to be extracted from partial rotation and topological entanglement entropy in Sec. III.2.1. These span all the obstructions to $\mathrm{U}(1)$-preserving gapped edge state mentioned above.
@@ VI Lieb-Schultz-Mattis type constraint
Suppose that we observed $\{\mathcal{T}^{f}_{1}(2\pi/p_{j},\pi/p_{j})\}$ has a non-trivial phase for a set of distinct prime numbers $\{p_{j}\}$. One can see that this leaves us two possibilities: 1. the edge cannot be gapped while preserving $\mathrm{U}(1)^{f}$ symmetry, or 2. the edge is gappable preserving $\mathrm{U}(1)^{f}$ symmetry, where the bulk must have multiple ground states on a torus with the degeneracy non-trivially bounded from below. This statement is reminiscent of the Lieb-Schultz-Mattis type theorems Lieb et al. (1961); Oshikawa (2000); Hastings (2004), which constrains the low-energy spectrum for a given input of the symmetry action on the ground state. Concretely, one can show the following theorem:
Theorem. For a (2+1)D topological order with $\mathrm{U}(1)^{f}$ symmetry, suppose that the partial rotation $\{\mathcal{T}^{f}_{1}(2\pi/p_{j},\pi/p_{j})\}$ has a non-trivial phase for a set of distinct prime numbers $\{p_{j}\}$. Then, the topological order admits a $\mathrm{U}(1)^{f}$ symmetry-preserving gapped edge state only if the number of anyons $r$ satisfies $r\geq r_{0}$, where $r_{0}$ is the smallest integer satisfying $\prod_{j}p_{j}\leq 2^{\frac{4}{3}r_{0}+8}3^{\frac{4}{3}r_{0}}$.
Proof. Let us assume that the topological order admits a $\mathrm{U}(1)^{f}$-preserving gapped edge, while the phases of $\{\mathcal{T}_{1}^{f}(2\pi/p_{j},2\pi/p_{j})\}$ are non-trivial. $\mathcal{T}^{f}_{1}(2\pi/p_{j},\pi/p_{j})$ obstructs the $\mathrm{U}(1)^{f}$-preserving gapped edge when $\gcd({N_{\mathrm{FS}}}(\mathcal{C}),p_{j})=1$, where $\mathcal{C}$ is the super-modular tensor category that describes the anyons of the topological order. We hence must have $\gcd({N_{\mathrm{FS}}}(\mathcal{C}),p_{j})>1$. This implies that ${N_{\mathrm{FS}}}(\mathcal{C})$ must be divisible by $\prod_{j}p_{j}$, so we have $\prod_{j}p_{j}\leq{N_{\mathrm{FS}}}(\mathcal{C})$. Meanwhile, let us consider a modular tensor category $\breve{\mathcal{C}}$ obtained by gauging ${\mathbb{Z}}_{2}^{f}$ symmetry of the fermionic topological phase $\mathcal{C}$. Mathematically, $\breve{\mathcal{C}}$ is the minimal modular extension of the super-modular category $\mathcal{C}$. For a modular category, it is known that ${N_{\mathrm{FS}}}(\breve{\mathcal{C}})\leq 2^{\frac{2\breve{r}}{3}+8}3^{\frac{2}{3}\breve{r}}$, where $\breve{r}$ is the number of anyons in $\breve{\mathcal{C}}$ Bruillard et al. (2015). In the presence of $\mathrm{U}(1)^{f}$ symmetry, there is an Abelian vison $v$ that carries ${\mathbb{Z}}_{2}^{f}$ vortex. In that case, one can write $\breve{r}=2r$, where $r$ is the number of anyons in $\mathcal{C}$. Using ${N_{\mathrm{FS}}}(\mathcal{C})\leq{N_{\mathrm{FS}}}(\breve{\mathcal{C}})$, we have
@@ VII.2 Review: anomaly and Hall conductivity of topological order with Lie group symmetry
$\begin{split}&\sum_{j}2\pi h(\pi_{\mathcal{A}}({\bf k}_{j}))\int_{W^{4}}\mathcal{P}(\pi_{\mathcal{A}}(w_{2}^{j}))\\ &+\sum_{j<k}2\pi m(\pi_{\mathcal{A}}({\bf k}_{j}),\pi_{\mathcal{A}}({\bf k}_{k}))\int_{W^{4}}\pi_{\mathcal{A}}(w_{2}^{j})\cup\pi_{\mathcal{A}}(w_{2}^{k})~{},\end{split}$ (41)
@@
$\sum_{j}\int_{C}\pi_{f}(g_{b}^{\ast}w_{2}^{j})=\int_{C}w_{2}(TW^{4})\quad\bmod 2,$ (42)
for any 2-cycle $C$. Here $g_{b}:W^{4}\to BG_{b}$ is the background $G_{b}$ gauge field, and $g_{b}^{*}w_{2}^{j}\in H^{2}(W^{4},{\mathbb{Z}}_{n_{j}})$ denotes the pull back of $w_{2}^{j}$. $w_{2}(TW^{4})$ is the second Stiefel-Whitney class of the tangent bundle of the spacetime $W^{4}$ on which the response action is supported.
@@ VII.4 Example: $G_{f}=\mathrm{SU}(2)^{f}$
$\frac{2\pi k}{4}\int_{W^{4}}\mathcal{P}(w_{2}(\mathrm{SO}(3))).$ (45)
This response action of $\mathrm{SU}(2)^{f}$ symmetry actually corresponds to a trivial SPT, when the Dirac quantization condition (42) is understood; $w_{2}(\mathrm{SO}(3))=w_{2}(TW^{4})$. In that case we have $\int\mathcal{P}(w_{2}(TW^{4}))=\sigma(W^{4})$ mod 4, with $\sigma(W^{4})$ the signature of a 4-manifold $W^{4}$. Hence, the above response action (45) is identified as the continuous gravitational theta angle, and topological order does not carry ’t Hooft anomaly of SU(2)f symmetry.
@@ VIII Discussions
Also, recent development of non-invertible symmetries uncovered the presence of various invariants associated with the responses of non-invertible symmetries Thorngren and Wang (2019); Zhang and Córdova (2023); Cordova et al. (2023). In particular, the electric Hall conductance can be defined for certain class of continuous non-invertible symmetries, which again forbids gapless edge modes preserving the symmetry Hsin et al. (2024). It would be interesting to extract such invariants from a microscopic wave function, as well as find higher versions of Hall conductivity for non-invertible symmetries.
Finally, the partial operations also allow us to characterize the invariants of crystalline topological phases with point group symmetry Shiozaki et al. (2017); Zhang et al. (2023); Kobayashi et al. (2024c). It would be interesting to develop a theory for the crystalline invariants of topological order, which is relevant to various invariants characterizing the crystalline spin liquids.
@@ A.1 Fermionic case
in terms of edge CFT with $\mathrm{U}(1)^{f}$ symmetry by utilizing the cut-and-glue approach Qi et al. (2012). For simplicity, let us consider the case where the edge CFT is described in terms of chiral rational CFT; the generalization to non-chiral case is rather straightforward, see Ref. Kobayashi et al., 2024b for the discussions in bosonic non-chiral phases.
@@
where we introduced the velocity $v$, correlation length $\xi_{l}=v\beta_{l}$, $\xi_{r}=v\beta_{r}$, and the circumference of the cylinder $L$. ${P}_{l}$ and ${P}_{r}$ are translation operators on the left and right edge ${P}_{l}=-\frac{1}{v}H_{l}$, ${P}_{r}=\frac{1}{v}H_{r}$. $\mathrm{Tr}_{\alpha}[]$ denotes the trace within the sector of Hilbert space on a circle labeled by the anyon $\alpha$. $\chi_{\alpha}(\tau;[s,\theta],[s^{\prime},\theta])$ with $s,s^{\prime}\in\{\mathrm{AP},\mathrm{P}\}$, $\theta,\theta^{\prime}\in[0,\pi)$ is the CFT character that corresponds to the partition function on a torus with modular parameter $\tau$ equipped with Spinc structure. For example,
@@
where $\alpha$ labels the topological sector within the anti-periodic sector, with the $\mathrm{U}(1)$ twisted boundary condition by $\theta$ in the spatial cycle. $\{L_{j}\}$ denotes the generator of the Virasoro algebra. In our setup where $L\ll\xi_{l}$, the CFT characters for the left edge are approximated as
@@ A.1.1 The case where $n$ is odd
$\begin{split}\chi_{a}\left(\frac{i\xi_{r}}{L}-\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}]\right)&=\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}S_{ab}\chi_{b}\left(-\frac{1}{\frac{i\xi_{r}}{L}-\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\\ &=\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}(ST^{n})_{ab}\chi_{b}\left(\frac{-in\frac{\xi_{r}}{L}}{\frac{i\xi_{r}}{L}-\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\end{split}$ (51)
@@
Here, $[]_{\pi}$ denotes the mod $\pi$ operation and the spin structure is acted by ${\mathbb{Z}}_{2}$ action as $\mathrm{AP}+1=\mathrm{P}$, $\mathrm{P}+1=\mathrm{AP}$.
[…]
$\mathcal{C}_{G_{f}}:=\left(\bigoplus_{g_{b}\in G_{b}}\mathcal{C}_{\mathrm{AP},g_{b}}\right)\oplus\left(\bigoplus_{g_{b}\in G_{b}}\mathcal{C}_{\mathrm{P},g_{b}}\right)$ (53)
The sum $b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}$ in Eq. (51) is sum over the simple objects in the sector $\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}$.
To evaluate the action of $S,T$ modular matrices, we note that the modular transformation in the twisted Hilbert space of CFT can be read by $S,T$ matrices of $G_{b}$-crossed braided fusion category $\mathcal{C}_{G_{f}}$ that describes the topological order with global symmetry in the bulk Barkeshli et al. (2019). In our case, the bosonic symmetry group is $G_{b}=\mathbb{R}/(\pi\mathbb{Z})$, and the relevant subgroup for the operation $U_{\frac{\pi}{n}}$ is the ${\mathbb{Z}}_{n}$ subgroup of $G_{b}$. Using the data of ${\mathbb{Z}}_{n}$-crossed braided fusion category, the $T$-matrix element is expressed as Barkeshli et al. (2019)
@@
$\begin{split}\chi_{a}\left(\frac{i\xi_{r}}{L}-\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}]\right)&=\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}(ST^{n})_{ab}\chi_{b}\left(\frac{-in\frac{\xi_{r}}{L}}{\frac{i\xi_{r}}{L}-\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\\ &=e^{-\frac{2\pi in}{24}c_{-}}\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}S_{ab}\theta_{b}^{n}\prod_{j=0}^{n-1}\eta_{b}\left(\frac{\pi}{n},\frac{j\pi}{n}\right)\cdot\chi_{b}\left(\frac{-in\frac{\xi_{r}}{L}}{\frac{i\xi_{r}}{L}-\frac{1}{n}};[\mathrm{AP},\frac{\pi}{n}],[\mathrm{AP},0]\right)\\ &=e^{-\frac{2\pi in}{24}c_{-}}\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}\sum_{c\in\mathcal{C}_{\mathrm{AP},0}}S_{ab}\theta_{b}^{n}\prod_{j=0}^{n-1}\eta_{b}\left(\frac{\pi}{n},\frac{j\pi}{n}\right)\cdot S_{bc}\chi_{c}\left(\frac{iL}{n^{2}\xi_{r}}+\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}]\right)\end{split}$ (55)
@@
$\mathcal{T}^{f}_{a}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)=e^{\frac{2\pi i}{n}(h_{a}-\frac{c_{-}}{24})}\frac{e^{-\frac{2\pi in}{24}c_{-}}\sum_{b\in\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}}\sum_{c\in\mathcal{C}_{\mathrm{AP},0}}(S_{a,b}+S_{a\psi,b})\theta_{b}^{n}\prod_{j=0}^{n-1}\eta_{b}\left(\frac{\pi}{n},\frac{j\pi}{n}\right)\cdot S_{bc}\chi_{c}(\frac{iL}{n^{2}\xi_{r}}+\frac{1}{n};[\mathrm{AP},0],[\mathrm{AP},\frac{\pi}{n}])}{\sum_{b\in\mathcal{C}_{\mathrm{AP},0}}(S_{a,b}+S_{a\psi,b})\chi_{b}(\frac{iL}{\xi_{r}};[\mathrm{AP},0],[\mathrm{AP},0])}$ (56)
@@
where $\mathcal{C}_{\mathrm{AP},\frac{\pi}{n}}$ is the twisted sector in $G_{b}$-crossed extension of super-modular category whose simple objects corresponds to the defects carrying $\pi/n\in G_{b}$. To relate this quantity to $\zeta^{H}_{n}$, we express the formula in terms of the anyons in the untwisted sector, instead of the twisted sector by $\frac{\pi}{n}$. This is done by using the consistency equations and the Verlinde formula of $G$-crossed braided fusion category given by Barkeshli et al. (2019)
@@
where we labeled ${\bf g}=\pi z_{1},{\bf h}=\pi z_{2},{\bf k}=\pi z_{3}$ with $z_{1},z_{2},z_{3}\in[0,1)$, and $v\in\mathcal{C}_{\mathrm{P,0}}$ is the vison defined as the relation $M_{b,v}=e^{i\pi Q_{b}}$ for all anyons $b\in\mathcal{C}_{\text{AP},0},\mathcal{C}_{\mathrm{P},0}$. The parameter $k\in{\mathbb{Z}}$ corresponds to stacking the bosonic IQH phase atop the state (here, the bosonic IQH is equivalent to eight copies of fermionic IQH state stacked with $\overline{E}_{8}$ state, so carries $\sigma_{H}=8$). The combination $h_{v}+k$ is identified with the electric Hall conductivity of the bosonic topological order as Manjunath and Barkeshli (2020)